OpenStudy (anonymous):

The multiplication property of equality could be used to solve which of the following equations?

OpenStudy (anonymous):

4x = 5 x + 5 = -4 = 1 (x + 5)(4x - 1) = 0 NEED HELP!!!!!!!!